OpenStudy (abmon98):

Use of the Data Booklet is relevant to this question. The approximate percentage composition of the atmosphere on four different planets is given in the table below. The density of a gas may be defined as the mass of 1 dm3 of the gas measured at s.t.p. Which mixture of gases has the greatest density? planet major gases / % by number of molecules A Jupiter H2 89.8, He 10.2 B Neptune H2 80.0, He 19.0, CH4 1.0 C Saturn H2 96.3, He 3.25, CH4 0.45 D Uranus H2 82.5, He 15.2, CH4 2.3

OpenStudy (aaronq):

Calculate the density for all the gases, assume that you have 1 L and use the molar masses of the gases, for example: For Uranus, the composition is 82.5% is \(H_2\) and 15.2% is He, 2.3% \(CH_4\) so the mass in 1 L \(= 82.5\%M_{H_2}+ 15.2\% M_{He}+2.3\%M_{CH_4}\) \(= 0.825(2g/mol)+ 0.152(2g/mol)+0.023(16g/mol)\) =2.322 g The density is the same because were using 1 L, if you need to show the work, then: \(\rho_{Uranus}=\dfrac{2.322g}{1L}=2.322~g/L\) (i'm assuming the percent composition by volume by way).

OpenStudy (abmon98):

how is helium 2g/mol if helium must have 4g/mol

OpenStudy (aaronq):

oh damn, sorry, i made a mistake.

OpenStudy (abmon98):

its okay but how did you reach to the number 2.322 did you divide the percentage by its mass

OpenStudy (aaronq):

multiplied the percentage of each gas by it's molar mass

OpenStudy (aaronq):

so it should be \(\rho_{Uranus}=2.626 ~g/L\)

OpenStudy (abmon98):

thank you so much ::) :D

OpenStudy (aaronq):

no problem! (thats not the final answer though)

OpenStudy (abmon98):

i know the idea know

OpenStudy (abmon98):

and i can solve the rest

OpenStudy (aaronq):

good stuff, i was just making sure.